OpenStudy (anonymous):

Solve (x^2 - 16)(x - 3)^2 + 9x^2

OpenStudy (anonymous):

@AQ99

OpenStudy (anonymous):

what is X?

OpenStudy (anonymous):

@563blackghost

OpenStudy (anonymous):

It doesnt say @Mycat4556

OpenStudy (anonymous):

k

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

does it say whats its equal to?

OpenStudy (anonymous):

This equation doesnt even have a rational solution.. >.< the roots are all irrational

OpenStudy (anonymous):

and no

OpenStudy (anonymous):

k then I dont know how to solve this... sorry

OpenStudy (anonymous):

thanks anyways

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

@AQ99

OpenStudy (anonymous):

@geny55

OpenStudy (aq99):

Um...Er...Im at a loss here...sorry :/

OpenStudy (anonymous):

Here's the partial solution: (x^2 - 16)(x - 3)^2 + 9x^2 = 0 x^2(x - 3)^2 - 16(x - 3)^2 + 9x^2 = 0 x^2(x^2 - 6x + 9) - 16(x - 3)^2 + 9x^2 = 0 x^4 - 6x^3 + 18x^2 - 16(x - 3)^2 = 0 x^4 - 6x^2(x - 3) - 16(x - 3)^2 = 0

OpenStudy (anonymous):

so confused ;.;

OpenStudy (anonymous):

i think thats right o.o idk this sucks -.-

OpenStudy (anonymous):

please someone help ill give medal and fan :/

OpenStudy (aq99):

@Hero

OpenStudy (triciaal):

to solve you need an equation can you verify the original example should +9x^2 be =9x^2?

OpenStudy (anonymous):

it just says solve @triciaal

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

just gonna close the question and quess at it

Nnesha (nnesha):

so it's + 9x^2 or = 9x^2 ?

Nnesha (nnesha):

anyways (x-3)^2 is same as (x-3) (x-3) apply foil method!

OpenStudy (triciaal):